OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. I have to find what Nakim did wrong and then simplify the expression. @phi

OpenStudy (anonymous):

\[3 \sqrt{2x}+ x \sqrt{8x}- 5\sqrt{18x}\]

OpenStudy (anonymous):

he got \[-10x \sqrt{2x}\]

OpenStudy (phi):

the first step is pull out any perfect squares from out of the square roots can you do that ?

OpenStudy (anonymous):

sqrt 2x

OpenStudy (phi):

? you can't simplify the first term, but you can simplify the 2nd and 3rd terms what do you get ?

OpenStudy (anonymous):

? 2nd term: sqrt 8x^2 I don't know what to do for the third. Is that what you mean?

OpenStudy (anonymous):

\[3 \sqrt2x + x \sqrt8x−5 \sqrt18x\] you can see that you have the eqaution so you can take out the second term by pulling out the perfect squares like this \[x \sqrt 8x\] turns into \[2x \sqrt2x\]

OpenStudy (anonymous):

can you do the third term?

OpenStudy (phi):

for the 2nd term you start with \[ x \sqrt{8x} \] break up the 8 into its prime factors. in this case 8= 2*2*2 \[ x \sqrt{2\cdot 2\cdot 2\cdot x} \] now "pull out" a pair of 2's, and put one of them out front \[ 2x \sqrt{ 2 x} \]

OpenStudy (anonymous):

Ohhh, I forgot about that. Okay. Second term is 2x sqrt 2x third is -45x sqrt 9x?

OpenStudy (phi):

how did you do the last term? First, notice that there is no x "out front" so how did an x show up there ? also, -5 turned into -45 which means you multiplied by 9. why ? you should do this step by step: factor the 18 into 2*3*3 you now have \[ -5\sqrt{2 \cdot 3\cdot 3\cdot x} \] now you can "pull out" *pairs* .... there is only one pair

OpenStudy (anonymous):

3 and 3, yes.

OpenStudy (phi):

so that means you pull out the pair of 3's (and only the 3's) and put one of them "out front"

OpenStudy (anonymous):

I see what I did wrong :P So, -15 sqrt 6x ?

OpenStudy (phi):

almost, you pull out a "pair" \[-5\sqrt{2 \cdot 3\cdot 3\cdot x} = -5\cdot 3\sqrt{2 \cdot \cancel{3\cdot 3}\cdot x}\]

OpenStudy (phi):

what do you have so far with all 3 terms ?

OpenStudy (anonymous):

Oh, so you don't keep one of the 3's behind at all?

OpenStudy (anonymous):

This is what I have so far http://www.sketchtoy.com/47619421

OpenStudy (phi):

looks good so far. as for how square roots work, you know \[ \sqrt{4} =2 \] but you can write that as \[ \sqrt{2\cdot 2} = 2 \] as a way to remember that you pull out a pair

OpenStudy (phi):

now factor out the \(\sqrt{2x} \) from each term

OpenStudy (anonymous):

Pull out the pairs?

OpenStudy (phi):

factoring is different, if you have (1A +xA +yA) for example, you can "factor out" the A to get (1+x+y)*A notice if you "distribute the A" you get: (A+Ax+Ay) so factoring is the opposite of distributing in your problem, factor out the \( \sqrt{2x}\)

OpenStudy (anonymous):

Make it all multiplied by 2?

OpenStudy (phi):

you have \[ 3 \sqrt{2x} +2x\sqrt{2x}-15\sqrt{2x}\] factor out \(\sqrt{2x}\) from each term. that means "erase it" from each term, put parens around everything, and put the \(\sqrt{2x} \)on the outside of the parens

OpenStudy (anonymous):

so multiplied by sqrt 2x ?

OpenStudy (phi):

try it. what do you get ?

OpenStudy (anonymous):

3+2x-15 * sqrt2x

OpenStudy (phi):

yes, except put the first 3 terms in parens (because the sqrt2x is multiplying all of them) (3+2x-15 )* sqrt2x now you have in the parens 3+2x+ -15 or (changing the order of adding -- which is allowed) 3+ -15 +2x can you simplify that ?

OpenStudy (anonymous):

-10x *sqrt 2

OpenStudy (phi):

that looks dubious, as Nakim got the same answer, and we know he is wrong

OpenStudy (phi):

you can combine *like terms* so you can add and subtract the pure numbers, but the 2x is not a pure number. Leave the 2x alone

OpenStudy (anonymous):

2x-12 *sqrt2

OpenStudy (phi):

remember the parens and what happen to sqrt(2x) (not just sqrt2)

OpenStudy (anonymous):

(2x-12)*sqrt2x

OpenStudy (phi):

yes. that could be the answer. you could also factor out a 2 from both terms can you do that ?

OpenStudy (anonymous):

No, I don't remember how.

OpenStudy (phi):

divide both terms by 2 and simplify then put a 2 outside the parens can you do that ?

OpenStudy (anonymous):

http://www.sketchtoy.com/47621827

OpenStudy (phi):

almost. First, you should not use x for multiply in algebra because it looks like the variable x. people either leave out the multiply sign or if you need it for clarity, use a dot \( \cdot \) meanwhile, you divided each term by 2, but you also have to put a 2 out front of the parens

OpenStudy (anonymous):

Can you show me? I'm a little confused

OpenStudy (phi):

example: (2x+4y)z factor out a 2 from each term: divide each term in the parens by 2 (x +2y)2z and put the 2 outside the parens

OpenStudy (anonymous):

Like this? http://www.sketchtoy.com/47622643

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Click the sketchtoy link http://www.sketchtoy.com/47622643

OpenStudy (ash2326):

Cool you see the second last step \[(2x-12)\times \sqrt {2x}\] You take 2 out of 2x-12 \[2(x-6)\times \sqrt{2x}\] Do you notice any diff here and the sketchy last line?

OpenStudy (anonymous):

That's all?

OpenStudy (ash2326):

yes, you missed a 2 in the square root. otherwise it's right

OpenStudy (anonymous):

Okay, can you help me with where Nakim messed up?

OpenStudy (ash2326):

that's very easy, if you think that there is an x out in all the terms then you'll get Nakim's answer. Hint lies in the fact that we have \[2x-12\] now if we have x along with 12, you'd get -10x, of course the square root of (2x) is there

OpenStudy (anonymous):

I don't understand...

OpenStudy (ash2326):

check this out http://www.sketchtoy.com/47622643

OpenStudy (ash2326):

oops thats yours

OpenStudy (ash2326):

ok I'll type here \[3 \sqrt 2x+x\sqrt{8x} - 5 \sqrt{18x}\] He'd have thought that x is out side every term \[3\underline x\sqrt{2x}+x\sqrt{8x}-5\underline x\sqrt{18x}\] \[3x\sqrt{2x}+2x\sqrt{8x}-15x\sqrt{2x}\] \[(3x+2x-15x)\sqrt{2x}\] \[-10x\sqrt{2x}\] Do you get this?

OpenStudy (anonymous):

Yes, Thanks!

OpenStudy (ash2326):

Are you sure? and you understood how it has to be done correctly?